OpenStudy (anonymous):

the product of two consecutive positive integers is 29 more than their sum. find the integers.

OpenStudy (anonymous):

\[n(n+1)=n+n+1+29\] is a start

myininaya (myininaya):

let x be the first integer the very next integer is x+1 x(x+1)=29+x+(x+1)